A fastball pitcher can throw a baseball at 100 km/h. If the windup and throw take 1.5 s what is the acceleration of the ball?
1
Expert's answer
2019-02-28T11:39:08-0500

The acceleration of the ball:

a=v/t

a=(100 km/h)/1.5s =(100 km/h∙(1000 m)/km∙1h/3600s )/1.5s=18.5 m/s^2